# Imbalance-Robust and Sampling-Efficient Continuous Conditional GANs via Adaptive Vicinity and Auxiliary Regularization

### v2

@@ 3.3 Auxiliary Generator Regularization
We propose a multi-task discriminator framework $D(\bm{x},y)$, with architecture illustrated in Figure 3 (a). In our implementation, the backbone of $D$ consists of the convolutional blocks from SNGAN (Miyato et al. 2018), denoted as $D_{\text{conv}}$, though this component can be readily replaced with alternative GAN architectures such as DCGAN (Radford, Metz, and Chintala 2015), SAGAN (Zhang et al. 2019), BigGAN (Brock, Donahue, and Simonyan 2019), or GigaGAN (Kang et al. 2023). The convolutional backbone is followed by three parallel fully-connected branches: an one-layer adversarial branch $f_{\text{adv}}$, a two-layer regression branch $f_{\text{reg}}$, and a three-layer density ratio estimation branch $f_{\text{dre}}$.
The adversarial output, denoted by $D_{\text{adv}}(\bm{x},y)$, is generated by processing the extracted features $\bm{h}^{\bm{x}}=D_{\text{conv}}(\bm{x},y)$ through the adversarial branch $f_{\text{adv}}$ combined with label projection (Miyato et al. 2018). This output is then used to compute the adversarial loss $\mathcal{L}^{D}_{\text{adv}}$ according to Eqs. (1) and (4). Notably, the proposed soft or hybrid AV mechanism is fully compatible with this computation.
The regression branch $f_{\text{reg}}$ predicts the label of the input image through $\hat{y}=f_{\text{reg}}(D_{\text{conv}}(\bm{x},y))$. To optimize this branch, we employ a $\gamma$-insensitive hinge loss adapted from support vector regression (Murphy 2022), formulated as:
@@
where $y_{i}^{r}$ and $y_{i}^{g}$ denote the ground truth label for $\bm{x}_{i}^{r}$ and the conditional label used to generate $\bm{x}_{i}^{g}$, respectively, while $\hat{y}_{i}^{r}$ and $\hat{y}_{i}^{g}$ represent the corresponding prediction from $f_{\text{reg}}$. The elasticity parameter $\gamma\geq 0$, empirically set as the maximum raidus $\kappa_{y_{i}^{r}+\epsilon}$ in a training batch, accounts for Gaussian noise $\epsilon$ in real labels and helps mitigate misalignment between $\bm{x}_{i}^{g}$ and its conditional label $y_{i}^{g}$. In practice, we often replace $y_{i}^{g}$ with $\tilde{y}^{g}_{i}$, where $\tilde{y}^{g}_{i}$ denotes the output of a ResNet-18 (He et al. 2016) separately trained for label regression on $\bm{x}_{i}^{g}$.
The DRE branch $f_{\text{dre}}$ predicts the conditional density ratio $r(\bm{x}|y)=p_{r}(\bm{x}|y)/p_{g}(\bm{x}|y)$ for real and fake images. This branch is optimized through the loss function $\mathcal{L}_{\text{dre}}^{D}$ (Eq. (5)), where $\lambda_{\text{dre}}=10^{-2}$. Unlike Ding et al. (2023a), our feature extractor $D_{\text{conv}}$ is jointly optimized with $f_{\text{dre}}$.
@@
where $\lambda^{D}_{\text{reg}}\geq 0$ and $\lambda^{D}_{\text{dre}}\geq 0$ are the weighting coefficients.
@@
This formulation directly minimizes the discrepancy between $\bm{x}_{i}^{g}$ and $\hat{y}_{i}^{g}$, thereby enforcing stronger label alignment during image synthesis. The $f$-divergence penalty, implemented through Pearson $\chi^{2}$ divergence with $f=(t-1)^{2}$ (Nowozin, Cseke, and Tomioka 2016), provides distributional matching:
@@
4 Experiments
4.1 Experimental Setup
@@
Datasets. Building upon established experimental protocols (Ding et al. 2023b; Ding, Wang, and Xu 2024; Ding et al. 2025), we conduct comprehensive evaluations across multiple benchmark datasets: RC-49 (Ding et al. 2023b) (44,051 chair images at $64\times 64$ resolution with rotation angles from $0.1^{\circ}$ to $89.9\circ$ in $0.1^{\circ}$ increments), UTKFace (Zhang, Song, and Qi 2017) (14,723 facial images across $64\times 64$, $128\times 128$, and $192\times 192$ resolutions with age labels from 1 to 60), and Steering Angle (Chen 2018) (12,271 driving images at $64\times 64$ and $128\times 128$ resolutions covering steering angles from $80^{\circ}$ to $80^{\circ}$). For UTKFace and Steering Angle experiments, all available samples are used for both training and evaluation. The RC-49 experiments employ a curated balanced subset of 11,250 images (25 samples each for 450 distinct angle labels) for training while utilizing the complete collection of 44,051 images for comprehensive evaluation.
@@
To rigorously evaluate model robustness to data imbalance, we introduce RC-49-I—a specialized variant of RC-49 comprising three subsets with distinct label distribution patterns (unimodal, bimodal, and trimodal; see Figure 5). This framework enables systematic comparison between three challenging imbalanced scenarios (UTKFace, Steering Angle, and RC-49-I) and the balanced RC-49 baseline. All models trained on RC-49-I are evaluated using the complete RC-49 dataset (44,051 samples) to ensure consistent assessment.
@@
Compared Methods. Our comparative evaluation examines six continuous conditional generative models: CcDPM (Zhao et al. 2024) and CCDM (Ding et al. 2025) as diffusion-based approaches; CcGAN (Ding et al. 2023b) and Dual-NDA (Ding, Wang, and Xu 2024) representing vanilla CcGAN architectures; along with our proposed variants CcGAN-AVAR-S (soft adaptive vicinity) and CcGAN-AVAR-H (hybrid adaptive vicinity).
Implementation Setup. We utilize the released results of CcDPM, CCDM, CcGAN, and Dual-NDA on RC-49, UTKFace, and Steering Angle from (Ding et al. 2025). For our newly introduced RC-49-I dataset, we implement these baseline methods while maintaining identical experimental configurations to those used for RC-49. As a key contribution of this work, we develop and release an optimized, unified codebase for CcGAN and our proposed CcGAN-AVAR framework, featuring support for multiple benchmark datasets, flexible backbone network selection, diverse vicinity selection strategies, various GAN losses and regularization penalties, exponential moving average (Karras et al. 2024), and mixed-precision training (Micikevicius et al. 2018).
@@
For CcGAN-AVAR-S and CcGAN-AVAR-H implementations, we set $N_{\text{AV}}$ to 50 (RC-49), 30 (RC-49-I), 400 (UTKFace), and 20 (Steering Angle). All experiments utilize a SNGAN backbone architecture with fixed hyperparameters $\lambda_{\text{reg}}^{D}=\lambda_{\text{reg}}^{G}=1$ and $\lambda_{f}^{G}=0.5$. The parameter $\lambda_{\text{dre}}^{D}$ is set to 1.0 for both variants on RC-49 and for CcGAN-AVAR-H on RC-49-I, while using 0.5 in all other cases. To compute the $\gamma$-insensitive hinge loss for the regression branch of $D$, we pre-train a ResNet-18 model on each dataset to predict labels from images. Complete training configurations are provided in the Appendix and our codebase.
@@
Evaluation Setup. For systematic evaluation, each trained model generates a predetermined number of images at specified evaluation centers across all datasets. On RC-49 and RC-49-I, models produce 200 fake images for each of 899 rotation angles spanning $0.1^{\circ}$ to $89.9$ in $0.1^{\circ}$ increments. For UTKFace, models generate 1,000 images per age from 1 to 60 years. The Steering Angle evaluation employs 2,000 uniformly distributed points across $80^{\circ},80^{\circ}]$, with 50 images synthesized per point. This protocol yields total generation counts of 179,800 images (RC-49 and RC-49-I), 60,000 images (UTKFace), and 100,000 images (Steering Angle) for each evaluated method.
@@
Following the evaluation protocol established in (Ding et al. 2023b; Ding, Wang, and Xu 2024; Ding et al. 2025), we assess generated images using one primary metric and three complementary metrics. The Sliding Fréchet Inception Distance (SFID) (Ding et al. 2023b) serves as our primary evaluation metric, computed as the averaged FID score across all evaluation centers. Additionally, we employ three complementary metrics: (1) the Naturalness Image Quality Evaluator (NIQE) (Mittal, Soundararajan, and Bovik 2012) for assessing visual realism, (2) Diversity (Ding et al. 2023b) for quantifying sample variety, and (3) Label Score (Ding et al. 2023b) for measuring label consistency. For SFID, NIQE, and Label Score, lower values indicate superior performance, whereas higher Diversity scores reflect greater sample variety. We refer interested readers to Ding et al. (2025) for complete technical details regarding these evaluation metrics.
@@
Furthermore, we evaluate the sampling efficiency of each compared model by reporting its sampling speed (images/second) and GPU memory consumption (GiB), measured on a Linux system equipped with an NVIDIA RTX 4090D.
4.2 Experimental Results
We conduct a comprehensive evaluation of candidate methods across nine distinct settings, including eight imbalanced configurations and one balanced setting. The empirical results, presented in Tables 1 and 2 along with Figure 6, reveal several key findings:

### v3

@@ III-C Auxiliary Generator Regularization
We propose a multi-task discriminator framework $D(\bm{x},y)$, with architecture illustrated in Fig. 3 (a) and detailed in Section III-D. In our implementation, the backbone of $D$ utilizes convolutional blocks from SNGAN [24], denoted as $D_{\text{conv}}$, which maps the input image $\bm{x}$ to extracted features $\bm{h}_{\bm{x}}=D_{\text{conv}}(\bm{x},y)$. This backbone is designed to be modular and can be readily replaced with other established GAN architectures such as DCGAN [25], SAGAN [26], or BigGAN [27]. The convolutional backbone is followed by three parallel linear branches: an one-layer adversarial branch $f_{\text{adv}}$, a two-layer regression branch $f_{\text{reg}}$, and a three-layer density ratio estimation branch $f_{\text{dre}}$.
The adversarial output, denoted by $D_{\text{adv}}(\bm{x},y)$, is generated by processing the extracted features $\bm{h}_{\bm{x}}$ through the adversarial branch $f_{\text{adv}}$ combined with label projection [24], i.e.,
$D_{\text{adv}}(\bm{x},y)=f_{\text{adv}}(\bm{h}_{\bm{x}})+\bm{h}_{\bm{x}}\cdot\bm{h}^{\prime}_{y},$ (13)
where $\bm{h}^{\prime}_{y}=f_{\text{proj}}(\bm{h}_{y})$ and $f_{\text{proj}}$ is a single linear layer ensuring dimensional consistency between $\bm{h}_{\bm{x}}$ and $\bm{h}^{\prime}_{y}$. The output $D_{\text{adv}}(\bm{x},y)$ is subsequently used to compute the adversarial loss $\mathcal{L}^{D}_{\text{adv}}$ according to Eq. (1) or its hinge loss variant, with sample weights given by Eq. (8) or Eq. (9)—corresponding to the SAVDL or HAVDL formulation, respectively.
The regression branch $f_{\text{reg}}$ predicts the label of the input image through $\hat{y}=f_{\text{reg}}(D_{\text{conv}}(\bm{x},y))$. To optimize this branch, we employ a $\gamma$-insensitive hinge loss adapted from support vector regression [28], formulated as:
@@
where $y_{i}^{r}$ and $y_{i}^{g}$ denote the ground truth label for $\bm{x}_{i}^{r}$ and the conditional label used to generate $\bm{x}_{i}^{g}$, respectively, while $\hat{y}_{i}^{r}$ and $\hat{y}_{i}^{g}$ represent the corresponding prediction from $f_{\text{reg}}$. The elasticity parameter $\gamma\geq 0$, empirically set as the maximum raidus $\kappa_{y_{i}^{r}+\epsilon}$ in a training batch, accounts for Gaussian noise $\epsilon$ in real labels and helps mitigate misalignment between $\bm{x}_{i}^{g}$ and its conditional label $y_{i}^{g}$. To address the inherent label inconsistency between $y_{i}^{g}$ and the actual ground truth label of $\bm{x}_{i}^{g}$, we substitute $y_{i}^{g}$ with $\tilde{y}^{g}_{i}$—the output of a separately trained ResNet-18 [29] regressor applied to $\bm{x}_{i}^{g}$.
The DRE branch $f_{\text{dre}}$ predicts the conditional density ratio $r(\bm{x}|y)=p_{r}(\bm{x}|y)/p_{g}(\bm{x}|y)$ for real and fake images through
$\hat{r}(\bm{x}|y)=f_{\text{dre}}(D_{\text{conv}}(\bm{x},y)).$ (15)
This branch is optimized by minimizing a penalized Softplus loss, $\mathcal{L}_{\text{dre}}^{D}$, which takes a form similar to Eq. (6):
$\mathcal{L}^{D}_{\text{dre}}=\frac{1}{N^{g}}\sum_{i=1}^{N^{g}}\left[\delta(\hat{r}(\bm{x}_{i}^{g}|y_{i}^{g}))\hat{r}(\bm{x}_{i}^{g}|y_{i}^{g})-\eta(\hat{r}(\bm{x}_{i}^{g}|y_{i}^{g}))\right]$
$-\frac{1}{N^{r}}\sum_{i=1}^{N_{r}}\delta(\hat{r}(\bm{x}_{i}^{r}|y_{i}^{r}))+\lambda_{\text{dre}}\cdot\left(\frac{1}{N^{g}}\sum_{i=1}^{N^{g}}\hat{r}(\bm{x}_{i}^{g}|y_{i}^{g})-1\right)^{2},$ (16)
where $\lambda_{\text{dre}}=10^{-2}$. Unlike [21], our feature extractor $D_{\text{conv}}$ is jointly optimized with $f_{\text{dre}}$ during minimization of $\mathcal{L}^{D}_{\text{dre}}$.
@@
where $\lambda^{D}_{\text{reg}}\geq 0$ and $\lambda^{D}_{\text{dre}}\geq 0$ are the weighting coefficients, and $\mathcal{L}^{D}_{\text{adv}}$ denotes SAVDL/HAVDL as defined in Section III-B.
@@
This formulation directly minimizes the discrepancy between $\bm{x}_{i}^{g}$ and $\hat{y}_{i}^{g}$, thereby enforcing stronger label consistency during image synthesis. The $f$-divergence penalty, implemented through Pearson $\chi^{2}$ divergence with $f=(t-1)^{2}$ [30], provides distributional matching:
@@
Remark 3 (Extra Illustration for $f$-Divergence Penalty).
The $f$-divergence [30] between $p_{r}(\bm{x}|y)$ and $p_{g}(\bm{x}|y)$ is defined as:
$=\mathbb{E}_{\bm{x}\sim p_{g}(\bm{x}|y)}\left[f\left(\frac{p_{r}(\bm{x}|y)}{p_{g}(\bm{x}|y)}\right)\right],$ (21)
$=\mathbb{E}_{\bm{x}\sim p_{g}(\bm{x}|y)}\left[(r(\bm{x}|y)-1)^{2}\right].$ (22)
Since $\hat{r}(\bm{x}|y)=f_{\text{dre}}(D_{\text{conv}}(\bm{x},y))$ in Eq. (15) is trained during the discriminator updates to estimate the conditional density ratio $r(\bm{x}|y)=p_{r}(\bm{x}|y)/p_{g}(\bm{x}|y)$, the $f$-divergence penalty defined in Eq. (19) approximates Eq. (22), and minimizing $\mathcal{L}^{G}_{f}$ encourages the generator to better match the target data distribution.
III-D Network Architecture Design
CcGAN-AVAR employs a SNGAN-based architecture [24] for both generator and discriminator networks, with a modified number of residual blocks in the generator compared to the original implementation in [16]. While Ding et al. [16] recommend SAGAN [26] for high-resolution scenarios, we observe limited compatibility between its self-attention mechanism and exponential moving average (EMA) [31]—a technique that effectively enhances SNGAN performance in our experiments.
A key contribution is our modified discriminator architecture, which integrates both a two-layer regression branch and a three-layer density ratio estimation (DRE) branch, as shown in Fig. 5. The regression branch uses two linear layers with spectral normalization (SN) [24] and ReLU activation, while the DRE branch replaces SN with group normalization (GN) [32] using 8 groups. In certain experimental settings, such as UTKFace ($\times192$), we introduce a dropout layer after the ReLU activation. This is not for conventional overfitting prevention, but to address a specific training issue: without it, the DRE branch fails to converge, as indicated by a stagnant training loss.
Figure 5: Architectures of the two-layer linear regression branch ($f_{\text{reg}}$) and the three-layer linear density ratio estimation (DRE) branch ($f_{\text{dre}}$) within the discriminator. Spectral normalization (SN) and group normalization (GN, 8 groups) are applied in their respective components. The inclusion of Dropout in $f_{\text{dre}}$ is optional. The integers within the linear layers indicate the number of output channels.
Remark 4 (Relation to ACGAN).
CcGAN-AVAR differs fundamentally from ACGAN [33, 34, 35] in three key aspects. First, while ACGAN uses an auxiliary classification branch for class-conditional generation, its discriminator does not directly receive class information. In contrast, CcGAN-AVAR integrates a regression branch for continuous label prediction and explicitly injects condition labels into the discriminator via label projection (Fig. 3). Second, whereas ACGAN relies on a cross-entropy loss for classification, CcGAN-AVAR employs a regression loss—specifically, a custom $\gamma$-insensitive hinge loss (Eq. (6)) designed to improve robustness against label noise, rather than a standard mean squared error. Third, and most distinctively, CcGAN-AVAR implements a density ratio estimation branch that imposes an $f$-divergence penalty during generator training—a feature entirely absent in standard ACGAN and other GAN architectures.
III-E A Unified and Scalable Codebase
Prior implementations of CcGANs provided by [15, 16, 23, 17, 21] suffer from several limitations: (1) separate codebases for each benchmark dataset rather than a unified framework; (2) inconvenient adaptation to custom datasets; (3) limited support for generator and discriminator architectures (only SNGAN [24] and SAGAN [26]); (4) lack of exponential moving average (EMA) [31] support; (5) no mixed-precision training [36], making them inefficient for GPUs with limited memory.
To address these limitations, we introduce a unified and scalable codebase with the following features:
  • A unified implementation supporting both the original CcGAN [16, 23] and the proposed CcGAN-AVAR.
  • A single-codebase solution for multiple benchmark datasets, designed for easy integration of new datasets via modifications to dataset.py.
  • Flexible network architecture support, including DCGAN [25], SNGAN [24], SAGAN [26], and BigGAN [27], with configurations automatically adapted for resolutions of $64\times 64$, $128\times 128$, $192\times 192$, and $256\times 256$. This flexibility allows us to investigate the impact of the architecture on performance. Our ablation study in Section IV reveals that the simpler SNGAN performs best, a finding that aligns with the design rationale outlined in Section III-D.
  • Multiple vicinity selection mechanisms: vanilla hard/soft vicinity [16, 23], and the proposed soft adaptive vicinity (SAV) and hybrid adaptive vicinity (HAV).
  • Training enhancements such as EMA for improved generation quality and mixed-precision training (via Accelerate [37]) for reduced memory usage.
IV Experiments
IV-A Experimental Setup
@@
Datasets. Building upon established experimental protocols [16, 23, 14], we conduct comprehensive evaluations across multiple benchmark datasets: RC-49 [16] (44,051 chair images at $64\times 64$ resolution with rotation angles from $0.1^{\circ}$ to $89.9^{\circ}$ in $0.1^{\circ}$ increments), UTKFace [38] (14,723 facial images across $64\times 64$, $128\times 128$, , and $192\times 192$ resolutions with age labels from 1 to 60), and Steering Angle [39] (12,271 driving images at $64\times 64$ and $128\times 128$ resolutions covering steering angles from $80^{\circ}$ to $80^{\circ}$). For UTKFace and Steering Angle experiments, all available samples are used for both training and evaluation. The RC-49 experiments employ a curated balanced subset of 11,250 images (25 samples each for 450 distinct angle labels) for training while utilizing the complete collection of 44,051 images for comprehensive evaluation.
@@
To rigorously evaluate model robustness to data imbalance, we introduce RC-49-I—a specialized variant of RC-49 comprising three subsets with distinct label distribution patterns (unimodal, bimodal, and trimodal; see Fig. 6). This framework enables systematic comparison between three challenging imbalanced scenarios (UTKFace, Steering Angle, and RC-49-I) and the balanced RC-49 baseline. All models trained on RC-49-I are evaluated using the complete RC-49 dataset (44,051 samples) to ensure consistent assessment.
We also create $256\times 256$ versions of UTKFace and Steering Angle from their lower-resolution counterparts via super-resolution methods [40, 41]. Detailed procedures and example images are provided in the Appendix.
@@
Compared Methods. Our comparative evaluation examines six continuous conditional generative models: CcDPM [2] and CCDM [14] as diffusion-based approaches; CcGAN [16] and Dual-NDA [23] representing vanilla CcGAN architectures; along with our proposed variants CcGAN-AVAR-S (soft adaptive vicinity) and CcGAN-AVAR-H (hybrid adaptive vicinity).
Implementation Setup. We utilize the released results of CcDPM, CCDM, CcGAN, and Dual-NDA on RC-49, UTKFace, and Steering Angle from [14]. For our newly introduced RC-49-I dataset, we implement these baseline methods while maintaining identical experimental configurations to those used for RC-49. For the $256\times 256$ versions of UTKFace and Steering Angle, we re-implement all methods based on their original lower-resolution setups but reduce the batch size to accommodate GPU memory constraints.
@@
We implement CcGAN-AVAR by first establishing a strong CcGAN baseline using our unified codebase, employing SNGAN as the backbone architecture with exponential moving average and mixed-precision training enabled. This baseline achieves performance comparable to or better than the implementation in [16], with results summarized in Table III of Section IV-C. Building on this strong baseline, we develop CcGAN-AVAR-S and CcGAN-AVAR-H, setting $N_{\text{AV}}$ to 50 (RC-49), 30 (RC-49-I), 400 (UTKFace), and 20 (Steering Angle), and fixing hyperparameters $\lambda_{\text{reg}}^{D}=\lambda_{\text{reg}}^{G}=1$ and $\lambda_{f}^{G}=0.5$. The parameter $\lambda_{\text{dre}}^{D}$ is set to 1.0 for both variants on RC-49 and for CcGAN-AVAR-H on RC-49-I, while using 0.5 in all other cases. To compute the $\gamma$-insensitive hinge loss for the regression branch of $D$, we pre-train a regression-oriented ResNet-18 model on real image-label pairs $\{(\bm{x}^{r}_{i},y^{r}_{i})\}_{i=1}^{N^{r}}$. Complete training configurations are provided in the Appendix and our codebase.
DiffAugment [42], an effective data augmentation strategy, is applied to all GAN-based methods. Additionally, all GAN-based methods employ hinge loss–based vicinal adversarial objectives (see Eq. (S.26) in the Appendix of [16]).
@@
Evaluation Setup. For systematic evaluation, each trained model generates a predetermined number of images at specified evaluation centers across all datasets. On RC-49 and RC-49-I, models produce 200 fake images for each of 899 rotation angles spanning $0.1^{\circ}$ to $89.9^{\circ}$ in $0.1^{\circ}$ increments. For UTKFace, models generate 1,000 images per age from 1 to 60 years. The Steering Angle evaluation employs 2,000 uniformly distributed points across $80^{\circ},80^{\circ}]$, with 50 images synthesized per point. This protocol yields total generation counts of 179,800 images (RC-49 and RC-49-I), 60,000 images (UTKFace), and 100,000 images (Steering Angle) for each evaluated method.
@@
Following the evaluation protocol established in [16, 23, 14], we assess generated images using one primary metric and three complementary metrics. The Sliding Fréchet Inception Distance (SFID) [16] serves as our primary evaluation metric, computed as the averaged FID score across all evaluation centers. Additionally, we employ three complementary metrics: (1) the Naturalness Image Quality Evaluator (NIQE) [43] for assessing visual realism, (2) Diversity [16] for quantifying sample variety, and (3) Label Score [16] for measuring label consistency. For SFID, NIQE, and Label Score, lower values indicate superior performance, whereas higher Diversity scores reflect greater sample variety. We refer interested readers to [14] for complete technical details regarding these evaluation metrics.
@@
Furthermore, we evaluate the sampling efficiency of each model under a fixed sampling batch size per setting to ensure comparability. The reported metrics—sampling speed (images/second) and GPU memory consumption (GiB)—were measured on a Linux system with NVIDIA RTX 4090D.
IV-B Experimental Results
We conduct a comprehensive evaluation across eleven distinct settings: ten imbalanced configurations (RC-49, RC-49-I, UTKFace, and Steering Angle from $64\times 64$ to $256\times 256$) and one balanced setting (RC-49). The empirical results, presented in Tables I and II together with Fig. 7, demonstrate that the proposed CcGAN-AVAR framework achieves an optimal balance between generation quality and sampling efficiency. Compared to GAN-based approaches, it delivers consistently superior generation performance across all experimental settings while maintaining comparable—or even better—sampling efficiency. When evaluated against diffusion-based baselines, CcGAN-AVAR achieves a 300× to 2000× acceleration in generation time while matching or exceeding their generation quality. Detailed findings are shown as follows: